# Optimal control under unknown intensity with Bayesian learning

### v1

@@ 2.2 Girsanov’s theorem and the posterior distribution
We observed that our optimization problem does not depend on the initial probability space. We propose here a specific and convenient construction of these objects, using Girsanov’s theorem.
We consider $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{Q})$ a probability space such that $(N_{t})$ is a standard $({\mathcal{F}}_{t})$-Poisson process with rate $1$. We denote by $(\tau_{k})$ the successive jump times of $(N_{t})$. We write ${\mathcal{F}}^{N}_{t}=\sigma(N_{r},r\leq t)$ for the natural filtration of $N$.
Let ${\mathcal{A}}$ be the collection of all the ${\mathcal{F}}^{N}$-predictable processes $\gamma$ such that
@@
Given $\gamma\in{\mathcal{A}}$ and $y\in\mathbb{R}$, we consider $(Y_{t})$ the solution of the SDE
@@
has a unique locally bounded solution ($\sup_{u\in[0,t]}|L_{u}|<\infty~{}a.s.$) and this solution is $L_{t}(\lambda)$ as defined above. In addition, $t\mapsto L_{t}(\lambda)$ is a $(\mathbb{Q},({\mathcal{F}}^{N}_{t}))$-local martingale.
@@
This follows from [5][Th. A.T4] ∎
Consider now $\Lambda$ a ${\mathcal{F}}_{0}$-random variable, with probability law $\mu$. As $(N_{t})$ is a $({\mathcal{F}}_{t})$ standard Poisson process, it holds that $\Lambda$ and $(N_{t})_{t\geq 0}$ are independent under $\mathbb{Q}$.
@@
Under Assumption 2.1, $\mathbb{E}_{\mathbb{Q}}L_{T}(\Lambda)=1$ and so $(L_{t}(\Lambda))$ is a $(\mathbb{Q},({\mathcal{F}}_{t}))$-martingale.
@@
By assumptions, for all $\lambda\in\mathbb{R}_{+}$,
@@
We first prove that $\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$ for all $\lambda\in\mathbb{R}_{+}$. We have $L_{t}\leq C(\lambda)^{N_{t}}e^{t}$ and $\lvert f_{\lambda}(Y_{t-})-1\rvert\leq C(\lambda)+1$. Therefore,
@@
By [5, T8, p. 27], this implies that $\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$. Finally note that there exists a measurable function $\Phi$ such that
@@
On $\mathbb{Q}$, $(N_{t})_{t\geq 0}$ and $\Lambda$ are independent. Therefore,
@@
where $\Psi(\lambda):=\mathbb{E}_{\mathbb{Q}}\Phi(N_{\cdot\wedge T},\lambda)=\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$. We deduce that $\mathbb{E}_{\mathbb{Q}}L_{T}(\Lambda)=\mathbb{E}_{\mathbb{Q}}\Psi(\Lambda)=1$. ∎
By Girsanov’s theorem, we deduce that
@@
On $\mathrm{d}\mathbb{P}:=L_{T}(\Lambda)\mathrm{d}\mathbb{Q}$, it holds that $(N_{t})$ is a point process of stochastic intensity $f_{\Lambda}(Y_{t-})$. For any measurable test function ϕ:+:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}_{+}$, consider:
@@
This follows from [5, Th. 3] and [5, Lem. 5]. For completeness, we give the arguments for the second part of the result. Take $A\in{\mathcal{F}}^{N}_{t}$. By definition of conditional expectation, we have $\mathbb{E}_{\mathbb{P}}[\mathbbm{1}_{A}\phi(\Lambda)]=\mathbb{E}_{\mathbb{P}}[\mathbbm{1}_{A}\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]]$. Therefore,
@@
This is holds for all $A\in{\mathcal{F}}^{N}_{t}$, we find that:
@@
giving the first equality for $\langle\phi,M^{\gamma}_{t}\rangle$. For the second equality, note that on $\mathbb{Q}$, $(N_{t})$ and $\Lambda$ are independent, therefore:
@@
Therefore, it holds that $a.s.$,
@@
Following the Bayesian terminology, $M^{\gamma}_{t}$ is the posterior distribution. In addition, the variance of the posterior distribution $V(N):=Var(M^{\gamma}_{T})=\mathbb{E}_{\mathbb{P}}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{T}]-\left(\mathbb{E}_{\mathbb{P}}(\Lambda~{}|~{}{\mathcal{F}}^{N}_{T})\right)^{2}$ is equal to:
@@
Let ϕ:+:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}_{+}$ be a non-negative measurable function. Then, $t\mapsto\langle\phi,M^{\gamma}_{t}\rangle=\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]$ is a martingale. In other words, $M^{\gamma}_{t}$ is a measure-valued martingale. In particular, it holds that $\mathbb{E}_{\mathbb{P}}\mathbb{E}_{\mathbb{P}}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{t}]=\mathbb{E}\Lambda^{2}<\infty$. Therefore, for any admissible control $\gamma$, we find that
@@
In addition, by the law of the total variance, we have:
@@
This shows in particular that, in expectation, the variance of $M^{\gamma}_{t}$ can only decay with time.
Altogether, we have proven that our optimization problem (2.1) is equal to:
@@
We again use the fact that under $\mathbb{Q}$, the standard Poisson process $N$ of rate $1$ and $\Lambda$ are independent.
@@ 2.3 The value function
We now slightly generalize the formulation of the problem to define the value function.
@@
We consider the canonical space $(\Omega,{\mathcal{F}},({\mathcal{F}}^{N}_{t}),\mathbb{Q})$ associated to a standard Poisson process $(N_{t})$ with rate $1$. We denote by $(\tau_{k})$ the successive jump times of $(N_{t})$.
@@
Denote by ${\mathcal{A}}$ the set of all $({\mathcal{F}}^{N}_{t})$-predictable processes $\gamma$ such that almost surely, $\int_{0}^{T}\gamma^{2}_{u}\mathrm{d}u<\infty$.
@@
Given $\gamma\in{\mathcal{A}},y\in\mathbb{R}$ and $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$, we let:
$J^{\gamma}(t,y,\mu):=\mathbb{E}_{\mathbb{Q}}\left[\int_{\mathbb{R}_{+}}\left(\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s+\text{Var}(M^{t,y,\mu,\gamma}_{T})\right)L^{y,\gamma}_{t,T}(\lambda)\mu(\mathrm{d}\lambda)\right],$
where for all $s\in[t,T]$:
@@
this equality being true for any non-negative measurable function ϕ:++:italic-$\phi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$. Moreover, the variance of $M^{t,y,\mu,\gamma}_{T}$ is:
@@
Finally, the value function is
$v(t,y,\mu):=\inf_{\gamma\in{\mathcal{A}}}J^{\gamma}(t,x,\mu).$
We have proven in the previous sections that:
@@ 2.4 Reduction to finite dimension
In what follows, we assume that:
@@
There is a globally Lipschitz and bounded function $g:\mathbb{R}\rightarrow\mathbb{R}_{+}$ such that:
@@
Surprisingly, with this assumption, the optimization problem over the space of probability measures can be reduced to a finite dimension problem. To proceed, given $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$, we set for all $z\geq 0$ and $n\in\mathbb{N}$:
$\tilde{v}(t,x,z,n):=v(t,x,m_{\mu}(n,z)),\quad\text{ where }\quad m_{\mu}(n,z)(\mathrm{d}\lambda):=\frac{\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda)}{\int_{\mathbb{R}_{+}}\theta^{n}e^{-\theta z}\mu(\mathrm{d}\theta)}.$
Note that $m_{\mu}(0,0)=\mu$. We also set
@@
Recall the definition of $M^{t,y,\mu,\gamma}_{s}$ and $Y^{t,y,\gamma}_{s}$ given by equations (2.4) and  (2.3). We define:
@@
Let $\gamma\in{\mathcal{A}}$. It holds that, almost surely:
@@
This follows from the fact that for every measurable non-negative function $\phi$, we have
@@
It holds that
$\tilde{v}(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[\int_{\mathbb{R}_{+}}\left(\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s+\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})\right)L^{t,y,\gamma}_{T}(\lambda)m_{\mu}(n,z)(\mathrm{d}\lambda)\right].$
The key point is that this optimization problem is now a classical stochastic optimization in finite dimension. For now on, we will concentrate our study on this finite-dimensional problem. Therefore, we write to simplify the notations $v(t,y,z,n)=\tilde{v}(t,y,z,n)$.
@@
Finally, we close this section with some properties satisfied by $m_{\mu}(n,z)$.
@@
Let $\Xi_{\mu}(n,z):=\Phi_{\mu}(n+1,z)/\Phi_{\mu}(n,z)$ be the expected value of the posterior distribution. We have for all $n\in\mathbb{N}$ and $z\geq 0$:
@@
The points 2 and 3 above show that the mean of posterior distribution decays between the jumps, while it always increase just after a jump. However, a similar behavior does not hold for the variance of the posterior distribution. Consider for instance the case where $\mu(\mathrm{d}\lambda)=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{\lambda_{\max}}$. After one jump, the posterior distribution is equal to $\delta_{\lambda_{\max}}$, with variance zero: $n\geq 1\implies\Psi_{\mu}(n,z)=0$. In that case, the variance decays to zero just after the first jump.
@@
Assume that $\text{Supp}(\mu)\subset[0,\lambda_{\max}]$. Then $\Psi_{\mu}(n,z)\leq\frac{\lambda^{2}_{\max}}{4}$ and
$\lvert\partial_{z}\Psi_{\mu}(n,z)\rvert\leq\frac{1}{2}\lambda^{3}_{\max}.$

### v2

@@ 2.3 Girsanov’s theorem and the posterior distribution
Having established that our optimization problem is independent of the initial probability space, we propose a specific and convenient construction of the relevant objects using Girsanov’s theorem.
Consider a probability space $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{Q})$ such that $(N_{t})$ is a standard $({\mathcal{F}}_{t})$-Poisson process with intensity $1$. Let $(\tau_{k})$ denote the successive jump times of $(N_{t})$, and let ${\mathcal{F}}^{N}_{t}=\sigma(N_{r},r\leq t)$ for the natural filtration of $N$.
Let ${\mathcal{A}}$ denote the set of all ${\mathcal{F}}^{N}$-predictable processes $\gamma$ such that
@@
Given $\gamma\in{\mathcal{A}}$ and $y\in\mathbb{R}$, we consider $(Y_{t})$ as the solution to the stochastic differential equation:
@@
has a unique locally bounded solution ($\sup_{u\in[0,t]}|L_{u}|<\infty~{}a.s.$), given by $L_{t}(\lambda)$ as defined above. Furthermore, $t\mapsto L_{t}(\lambda)$ is a $(\mathbb{Q},({\mathcal{F}}^{N}_{t}))$-local martingale.
@@
This follows from [6, Th. A.T4]. ∎
Now, consider $\Lambda$ an ${\mathcal{F}}_{0}$-measurable random variable with probability distribution $\mu$. Since $(N_{t})$ is a standard $({\mathcal{F}}_{t})$ Poisson process under $\mathbb{Q}$, $\Lambda$ and $(N_{t})_{t\geq 0}$ are independent under $\mathbb{Q}$.
@@
Under Assumption 2.1, $\mathbb{E}_{\mathbb{Q}}L_{T}(\Lambda)=1$, and thus $(L_{t}(\Lambda))$ is a $(\mathbb{Q},({\mathcal{F}}_{t}))$-martingale.
@@
By assumption, for all $\lambda\in\mathbb{R}_{+}$,
@@
We first show that $\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$ for all $\lambda\in\mathbb{R}_{+}$. Since $L_{t}\leq C(\lambda)^{N_{t}}e^{t}$ and $\lvert f_{\lambda}(Y_{t-})-1\rvert\leq C(\lambda)+1$, it follows that:
@@
By [6, T8, p. 27], this implies $\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$. Next, note that there exists a measurable function $\Phi$ such that:
@@
Since $(N_{t})_{t\geq 0}$ and $\Lambda$ are independent under $\mathbb{Q}$, we have:
@@
where $\Psi(\lambda):=\mathbb{E}_{\mathbb{Q}}\Phi(N_{\cdot\wedge T},\lambda)=\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$. Thus:
$\mathbb{E}_{\mathbb{Q}}L_{T}(\Lambda)=\mathbb{E}_{\mathbb{Q}}\Psi(\Lambda)=1.$
By Girsanov’s theorem, we obtain the following:
@@
On $\mathrm{d}\mathbb{P}:=L_{T}(\Lambda)\mathrm{d}\mathbb{Q}$, $(N_{t})$ is a point process with stochastic intensity $f_{\Lambda}(Y_{t-})$. For any measurable test function ϕ:+:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}_{+}$, define:
@@
This follows from [6, Th. 3] and [6, Lem. 5]. For completeness, we provide the arguments for the second part of the result. Let $A\in{\mathcal{F}}^{N}_{t}$. By definition of conditional expectation, we have $\mathbb{E}_{\mathbb{P}}[\mathbbm{1}_{A}\phi(\Lambda)]=\mathbb{E}_{\mathbb{P}}[\mathbbm{1}_{A}\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]]$. Thus,
@@
Since this holds for all $A\in{\mathcal{F}}^{N}_{t}$ it follows that:
@@
yielding the first equality for $\langle\phi,M^{\gamma}_{t}\rangle$. For the second equality, since $(N_{t})$ and $\Lambda$ are independent under $\mathbb{Q}$, we have::
@@
Thus, $a.s.$, the posterior distribution is given by:
@@
Additionally, the variance of the posterior distribution, $V(N):=Var(M^{\gamma}_{T})=\mathbb{E}_{\mathbb{P}}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{T}]-\left(\mathbb{E}_{\mathbb{P}}(\Lambda~{}|~{}{\mathcal{F}}^{N}_{T})\right)^{2}$, is given by:
@@
Let ϕ:+:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}_{+}$ be a non-negative measurable function. Then, $t\mapsto\langle\phi,M^{\gamma}_{t}\rangle=\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]$ is a martingale. In other words, $M^{\gamma}_{t}$ is a measure-valued martingale. In particular, it follows that $\mathbb{E}_{\mathbb{P}}\mathbb{E}_{\mathbb{P}}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{t}]=\mathbb{E}\Lambda^{2}<\infty$. Thus, for any admissible control $\gamma$, we have:
@@
Furthermore, by the law of total variance:
@@
This implies that, in expectation, the variance of $M^{\gamma}_{t}$ decreases over time.
Thus, we have shown that our optimization problem (2.2) can be expressed as:
@@
Here, we have again used the fact that under $\mathbb{Q}$, the standard Poisson process $N$ with intensity $1$ and $\Lambda$ are independent.
@@ 2.4 The value function
We now generalize the formulation of the problem to define the value function.
@@
We consider the canonical space $(\Omega,{\mathcal{F}},({\mathcal{F}}^{N}_{t}),\mathbb{Q})$ associated with a standard Poisson process $(N_{t})$ with intensity $1$. Specifically, $\Omega$ is the space of right-continuous simple counting functions, where $\Omega\ni\omega=(\omega_{t})_{t\geq 0}$ satisfies $\omega_{0}=0$ and is a piecewise constant function with a non-accumulating infinite sequence of jumps of size one. We define $N_{t}(\omega)=\omega_{t}$ for all $t\geq 0$. The space $\Omega$ is equipped with the Skorokhod topology, making it a Polish space. In addition, ${\mathcal{F}}$ is the associated Borelian $\sigma$-algebra, which is equal to $\sigma(N_{t},t\geq 0)$, and we define ${\mathcal{F}}^{N}_{t}=\sigma(N_{s},s\leq t)$. The measure $\mathbb{Q}$ is law of a standard Poisson process. We denote by $(\tau_{k}(\omega))$ the successive jump times of $(N_{t}(\omega))$, with the convention $\tau_{0}(\omega)=0$.
@@
Following [5], for $t\geq 0$, we define the translated filtration $({\mathcal{F}}^{N,t}_{s})_{s\geq 0}$ as:
${\mathcal{F}}^{N,t}_{s}:=\sigma(N_{r}-N_{t},r\in[t,t\vee s]).$
This filtration satisfies the property that, for all $t,s\in\mathbb{R}_{+}$, , a random variable measurable with respect to ${\mathcal{F}}^{N,t}_{s}$ is independent of ${\mathcal{F}}^{N}_{t}={\mathcal{F}}^{N,0}_{t}$. We denote by ${\mathcal{A}}_{t}$ the set of all $({\mathcal{F}}^{N,t}_{s})$-predictable processes $\gamma$ such that, almost surely, $\int_{0}^{T}\gamma^{2}_{u}\mathrm{d}u<\infty$.
@@
Given $\gamma\in{\mathcal{A}}_{t},y\in\mathbb{R}$, and $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$, we define:
$J^{\gamma}(t,y,\mu):=\mathbb{E}_{\mathbb{Q}}\left[\int_{\mathbb{R}_{+}}\left(\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s+\kappa\text{Var}(M^{t,y,\mu,\gamma}_{T})\right)L^{y,\gamma}_{t,T}(\lambda)\mu(\mathrm{d}\lambda)\right],$
where, for all $s\in[t,T]$:
@@
for any non-negative measurable function ϕ:++:italic-$\phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$. The variance of $M^{t,y,\mu,\gamma}_{T}$ is given by:
@@
The value function is defined as:
$v(t,y,\mu):=\inf_{\gamma\in{\mathcal{A}}_{t}}J^{\gamma}(t,y,\mu).$
We have shown in the previous sections that:
@@ 2.5 Reduction to finite dimension
In what follows, we assume the following:
@@
There is a globally Lipschitz and bounded function $g:\mathbb{R}\rightarrow\mathbb{R}_{+}$, such that:
@@
Under this assumption, the optimization problem over the space of probability measures can be reduced to a finite-dimensional problem. To proceed, given $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$, we define for all $z\geq 0$ and $n\in\mathbb{N}$:
$\tilde{v}(t,y,z,n):=v(t,y,m_{\mu}(n,z)),\quad\text{ where }\quad m_{\mu}(n,z)(\mathrm{d}\lambda):=\frac{\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda)}{\int_{\mathbb{R}_{+}}\theta^{n}e^{-\theta z}\mu(\mathrm{d}\theta)}.$
Note that $m_{\mu}(0,0)=\mu$. We also define:
@@
Recall the definition of $M^{t,y,\mu,\gamma}_{s}$ and $Y^{t,y,\gamma}_{s}$ given by equations (2.6) and  (2.5). We introduce:
@@
Let $\gamma\in{\mathcal{A}}_{t}$. It holds a.s. that:
@@
This follows from the fact that, for every non-negative measurable function $\phi$, we have:
@@
It holds that:
$\tilde{v}(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}_{t}}\mathbb{E}_{\mathbb{Q}}\left[\int_{\mathbb{R}_{+}}\left(\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s+\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})\right)L^{t,y,\gamma}_{T}(\lambda)m_{\mu}(n,z)(\mathrm{d}\lambda)\right].$
The key observation is that this optimization problem is now a classical finite-dimensional stochastic optimization problem. Henceforth, we focus our analysis on this finite-dimensional problem. To simplify notation, we write $v(t,y,z,n)=\tilde{v}(t,y,z,n)$.
@@
We conclude this section by establishing some properties of $m_{\mu}(n,z)$.
@@
Let $\Xi_{\mu}(n,z):=\Phi_{\mu}(n+1,z)/\Phi_{\mu}(n,z)$ denote the expected value of the posterior distribution. For all $n\in\mathbb{N}$ and $z\geq 0$, the following hold:
@@
Points 2 and 3 above indicate that the mean of the posterior distribution decreases between jumps but increases immediately after a jump. However, the variance of the posterior distribution does not exhibit a similar behavior. For example, consider the case where $\mu(\mathrm{d}\lambda)=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{\lambda_{\max}}$. After one jump, the posterior distribution becomes $\delta_{\lambda_{\max}}$, with zero variance: for $n\geq 1,\Psi_{\mu}(n,z)=0$. In this case, the variance drops to zero immediately after the first jump.
@@
Assume that $\text{Supp}(\mu)\subset[0,\lambda_{\max}]$. Then:
$\Psi_{\mu}(n,z)\leq\frac{\lambda^{2}_{\max}}{4}\ \text{ and }\ \lvert\partial_{z}\Psi_{\mu}(n,z)\rvert\leq\frac{1}{2}\lambda^{3}_{\max}.$